OpenStudy (anonymous):

PLEASE HELP :) The manager of a restaurant found that the cost of producing 100 cups of coffee is $11.02, while the cost of producing 400 cups is $40.12. Assume that the cost C(x) is a linear function of x, the number of cups produced. A. Find a formula for C(x)

jimthompson5910 (jim_thompson5910):

Let x = number of cups y = cost "the cost of producing 100 cups of coffee is $11.02" so we know x = 100 and y = 11.02. We have one point (100,11.02) we're also told "the cost of producing 400 cups is $40.12" this means x = 400, y = 40.12. The other point is (400, 40.12)

jimthompson5910 (jim_thompson5910):

What is the slope of the line through (100,11.02) and (400, 40.12) ?

OpenStudy (anonymous):

Is it -.023? That's what I calculated :/ @jim_thompson5910

OpenStudy (anonymous):

OOps made a mistake, I will redo that!

jimthompson5910 (jim_thompson5910):

that's incorrect

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

.097! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

that's the slope

jimthompson5910 (jim_thompson5910):

Plug in the slope m = 0.097 and the point (x,y) = (100,11.02) into y=mx+b and solve for b y = mx + b y = 0.097x + b 11.02 = 0.097*100 + b ... plug in x = 100 and y = 11.02 11.02 = 9.7 + b b = ???

OpenStudy (anonymous):

1.32! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

m = 0.097 b = 1.32

jimthompson5910 (jim_thompson5910):

y = mx+b y = 0.097x+b y = 0.097x+1.32

jimthompson5910 (jim_thompson5910):

Then you replace y with C(x) to turn it into a function y = 0.097x+1.32 C(x) = 0.097x+1.32

jimthompson5910 (jim_thompson5910):

so that's your cost function

jimthompson5910 (jim_thompson5910):

Plug in x = 100 and tell me what you get

OpenStudy (anonymous):

11.02 :) @jim_thompson5910

OpenStudy (anonymous):

Thank you so very much! You're very helpful and I appreciate it greatly :)

jimthompson5910 (jim_thompson5910):

you're welcome and I recommend you also check x = 400

OpenStudy (anonymous):

40.12 :) Thanks!

jimthompson5910 (jim_thompson5910):

you're welcome